# Visual Enhanced Depth Scaling for Multimodal Latent Reasoning

### 16 Apr 2026 v2

@@ 3.2 Spatially-Coherent Finer Visual Replay
As mentioned earlier, we empirically reveal the gradient disparities between visual and textual tokens throughout the learning dynamics. Specifically, visual tokens consistently exhibit substantially larger gradient norms and fluctuations compared to textual tokens, indicating that visual representations remain under-optimized despite their critical role in multimodal reasoning. Motivated by these insights, we introduce the visual replay module to reinforce the engagement of visual cues via salient region detection, while enhancing fine-grained spatially-coherent perception capabilities via self-distillation supervision at each reasoning step. Attention-Guided Region Focus. Several works [Li_2025_ICCV, DBLP:conf/naacl/ZhangQSYYXWGTY25] have demonstrated that LLMs exhibit fundamental visual grounding capabilities. To identify visually salient regions, we aggregate attention weights across all transformer layers and attention heads to obtain a consolidated spatial focus map. Specifically, given the $l$-th layer and the $h$-th attention head, we compute the mean attention map $\bar{\mathbf{A}}^{(t)}$ at reasoning step $t$:
@@
Spatially-Coherent Regularization. Although leveraging learned attention within Transformers provides explainable visual locations, it often suffers from limited spatial continuity due to the scattered nature of selected tokens and introduces noise associated with the attention sink phenomenon [DBLP:conf/iclr/XiaoTCHL24]. To mitigate these issues and enhance fine-grained perception without external annotations, we introduce self-distillation supervision. This mechanism involves cropping the visual regions exhibiting spatial coherence, re-encoding them, and supervising the visual latents with these high-fidelity features.
@@
where the density function $\mathcal{N}(r,c)$ counts the visited tokens falling within the window $\mathcal{N}(r,c)=\sum_{i\in\mathcal{I}^{(t)}}\mathbb{I}\big[r\leq r_{i}<r+W,c\leq c_{i}<c+W\big].$ Here, $(r_{i},c_{i})$ denotes the row-column position of token $i$ on the $G\times G$ grid, and $\mathbb{I}[\cdot]$ is the indicator function. This greedy selection ensures the cropped region captures a coherent visual context rather than scattered details. Then, the selected window is projected to pixel coordinates in the original image via mathematical transformation, cropped, and resized to the standard encoder input resolution using bilinear interpolation. We then re-encode this refined patch through the same visual encoder to obtain fine-grained representations $\{\mathbf{f}_{i}\}_{i=1}^{N_{v}^{\prime}}$. A global pooling operation $\mathrm{Pool}(\cdot)$ is applied to yield a robust reference token $\mathbf{u}^{\text{ref}}$:
$\mathbf{u}^{\text{ref}}=\mathrm{Pool}\big(\{\mathbf{f}_{i}\}_{i=1}^{N_{v}^{\prime}}\big).$ (8)
Finally, we align the coarse-grained global token $\mathbf{b}^{(t)}=\mathrm{Pool}(\mathbf{B}^{(t)})$ with this high-fidelity reference using a lightweight token super-resolution module $\mathcal{F}_{\text{SR}}:\mathbb{R}^{D}\to\mathbb{R}^{D}$. We minimize the reconstruction error as follows:
@@ 3.3 Routing Depth Scaling
While the visual replay mechanism significantly enhances fine-grained grounding by introducing spatially-coherent visual latents, existing methods typically allocate uniform computational budgets across all tokens during contextual refinement. Our empirical analysis reveals a fixed-depth optimization dilemma, demonstrating that token representations exhibit heterogeneous optimization complexities during latent training, which necessitates adaptive reasoning depths. To address this limitation without modifying the pretrained VLM architecture, while effectively leveraging its inherent knowledge, we introduce a lightweight router that dynamically allocates additional reasoning steps exclusively to critical tokens at each iteration. Router Network. Let the input token sequence at the $t$-th reasoning step be denoted as $\mathbf{U}^{(t)}=[\mathbf{X}\|\mathbf{B}^{(1)}\|\mathbf{z}^{(1)}\|\cdots\|\mathbf{B}^{(t)}\|\mathbf{z}^{(t)}]\in\mathbb{R}^{P^{(t)}\times d}$, where $P^{(t)}$ represents the total sequence length and $d$ is the hidden dimension. To facilitate the illustration of our depth scaling mechanism, we decompose the forward pass into layer-wise computations. Specifically, within the $l$-th transformer layer, we first compute the intermediate feature $\mathbf{H}^{(l,t)}=f(\mathbf{U}^{(t)})\in\mathbb{R}^{P^{(t)}\times d}$, where $f(\cdot)$ denotes the transformation of the $l$-th transformer layer in LLM. Subsequently, a lightweight router network computes a scalar importance score for each token based on its corresponding hidden representation:
@@
For notational brevity, we omit the layer index $l$ and reasoning step $t$ in the following formulation, where $d$ denotes the refinement depth. For instance, the score vector $\mathbf{s}^{(l,t)}$ is simplified to $\mathbf{s}^{(d)}$. The function $f_{i\in T_{\alpha}(\mathbf{s}^{(d)})}(\cdot)$ indicates that the attention operation is restricted to the token subset $T_{\alpha}(\mathbf{s}^{(d)})$. Here, $\mathbf{h}^{(0)}_{i}$ represents the initial hidden state of the $i$-th token after the first forward pass, and $\mathbf{m}^{(d)}\in\mathbb{R}^{P^{(t)}\times P^{(t)}}$ denotes the attention mask corresponding to refinement step $d$. The condition $i\in T_{\alpha}(\mathbf{s}^{(d)})$ functions as a binary gating mechanism, ensuring that only tokens with importance scores exceeding the threshold undergo additional computation.
The router network dynamically determines whether to apply depth scaling based on the current contextual representations in a data-driven manner. Tokens identified as critical are iteratively processed through the transformation function for $d$ additional refinement steps, while non-critical tokens retain their prior representations to preserve computational efficiency. Finally, we aggregate the depth-wise refined representation with step-aware positional encoding to form the final hidden state incrementally:
@@
This adaptive scaling strategy effectively allocates deeper reasoning pathways to critical tokens, enabling the model to capture complex visual contexts and facilitate more profound reasoning capabilities.
Table 1: Comparison of various multimodal reasoning baselines across three benchmarks. We selected three datasets featuring detailed reasoning chains for training and evaluate on their test split, respectively. Three metrics are reported: Accuracy (%), Average number of Autoregressive Steps (# AR steps), and Average Generation Time (AVG. Time). Evaluations were conducted on the M3CoT, ScienceQA, and GQA benchmarks using Qwen2-VL-7B and Chameleon-7B. denotes the reimplementation for the methods with the same configuration with ours. No-CoT notes that directly predicts answers without generating intermediate steps.
@@
Multimodal CoT [zhang2023multimodal] 42.5 106.3 3.10 58.3 83.9 2.44
CCoT [MitraCCoT] 44.1 177.2 5.31 63.8 164.0 5.23 51.2 76.4 7.21
ICoT [gao2024interleaved] 46.0 96.5 2.86 65.4 77.4 2.28
SCAFFOLD [lei2024scaffolding] 44.9 170.8 5.14 62.5 162.3 4.91 48.7 72.8 6.72
Chain-of-Focus [zhang2025chain] 64.3 185.7 2.63 91.2 162.3 2.09 61.8 128.6 3.01
@@
Multimodal CoT [zhang2023multimodal] 30.6 110.5 3.62 50.7 98.7 3.33
CCoT [MitraCCoT] 31.4 168.4 5.35 51.3 174.2 5.39 33.1 150.6 5.31
ICoT [gao2024interleaved] 32.3 110.9 5.43 53.4 92.4 4.62
SCAFFOLD [lei2024scaffolding] 31.1 194.3 6.12 47.5 160.6 6.03 32.8 156.0 4.17
Chain-of-Focus [zhang2025chain] 36.5 739.4 3.09 61.2 717.1 2.56 34.6 360.4 2.98
@@
GPT-4o [gpt4v] 68.70 72.00 68.00 64.70 63.80 30.39 31.80
Qwen2.5-VL-7B [bai2025qwen2] 65.67 65.31 56.57 68.25 53.60 59.70 68.20 25.60 37.50
LLaVA-OneVision [li2024llava] 9M 74.00 61.22 51.10 63.00 49.34 59.13 58.60 - -
InternVL3.5-8B [wang2025internvl3] 70K 57.67 69.78 56.15 59.38 54.81 53.33 71.60 28.30 -
@@
Multimodal CoT [zhang2023multimodal] 68.10 54.11 63.60 - - 57.90 56.40 21.80 35.60
CCoT [MitraCCoT] 69.00 68.95 64.90 - - 58.70 57.80 22.50 36.30
ICoT [gao2024interleaved] 69.30 70.27 65.50 - - 60.40 58.90 23.30 37.00
@@
PAPO [wang2025perception] 39K 68.67 54.11 57.52 68.12 52.66 45.80 67.53 - -
Vision-R1 [huang2025vision] 200K 72.67 68.95 63.83 75.12 52.71 62.67 52.40 - 40.20
VL-Rethinker [wang2025vl] 39K 72.67 70.27 71.08 63.50 55.55 63.20 72.80 29.30 -
DeepEyes [zheng2025deepeyes] 47K 70.00 69.08 62.57 69.12 51.08 58.73 70.10 26.60 -
@@
LVR [li2025lvr] 470K 64.00 47.39 65.19 53.62 53.60 57.93 - - -
Monet [wang2025monet] 125K 68.00 65.88 56.36 68.00 50.71 60.33 - - -
DMLR [liu2025reasoningminddynamicmultimodal] 70.10 - 65.80 - 56.92 60.27 59.10 24.40 38.80
Laser [li2025latent] 267K 72.00 70.05 67.72 72.50 - 60.10 - - -
@@
Table 2: Performance comparison of our method against baselines across four paradigms: Zero-Shot VLMs, Explicit CoT, Tool-Use & RL, and Latent Reasoning. Benchmarks are categorized into three domains: Visual Perception (MMVP, Seed-Bench-2-Plus, HallusionBench, HR-Bench), Compositional Reasoning (BLINK, MMStar), and Mathematical Reasoning (MathVista, MathVision, $M$-Math). Among latent reasoning approaches, the best results are highlighted in bold, and the second-best are underlined. Our method is built upon the Qwen2.5-VL-7B backbone.
Table 3: We perform comprehensive ablation studies to analyze the contribution of each design choice, including hyper-parameter configurations and architectural components. Experiments are conducted on three CoT benchmarks with multiple backbone models. To save the training overhead, all models are trained for 16 epochs with a balanced sampled subset.
@@
3.4 Training Procedure
Curriculum Latent Training. To mitigate annotation overhead and avoid the risk of human priors constraining model learning, we depart from prior approaches that rely on intermediate supervision for latent representations. Instead, we design a curriculum that facilitates the contextual and logical dependencies between implicit latents and explicit reasoning chains. In the initial stage, the model is trained with standard Chain-of-Thought (CoT) supervision, generating all reasoning steps explicitly to establish foundational reasoning capabilities. Subsequently, as training progresses, latent tokens are incrementally introduced. Specifically, one explicit reasoning step is progressively encapsulated into an informative $\left<\text{latent}\right$> token. Through this curriculum, each latent token progressively learns to ground the relevant contextual cues, effectively internalizing explicit reasoning chains into compact latent representations.
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda\frac{1}{T_{r}}\sum_{t=1}^{T_{r}}\mathcal{L}_{\text{recon}}^{(t)},$ (13)
where $\mathcal{L}_{\text{CE}}$ is the standard cross-entropy loss over the language modeling task,
$\mathcal{L}_{\text{CE}}=-\sum_{t=1}^{T_{r}+T_{a}}\log p_{\theta}\left(a_{t}|\mathbf{U}^{(t)},a_{<t}\right),$ (14)
where $\lambda$ is a hyperparameter balancing the primary task and the auxiliary reconstruction objective. During inference, only the LLM component is active, and the visual play module is disabled, ensuring no additional computational overhead at test time.
4 Experiments
We first conducted extensive evaluations on twelve diverse benchmarks against state-of-the-art competitors to validate the superiority of our method. Then, we provide comprehensive analysis through multi-faceted ablation studies and in-depth investigations to elucidate the underlying mechanisms of our approach.
4.1 Experimental Setup
Evaluation Benchmark.
We evaluate our method on three tasks across twelve benchmarks: (1) Mathematics Reasoning (MathVista [mathvista], MathVision [mathvision], MM-Math [mmmath]); (2) Vision-centric Reasoning (Hallusion-Bench [hallusionbench], MMVP [mmvp], Seed-Bench-2-Plus [seedbench2plus], HR-Bench [hrbench], GQA [gqa]); (3) Multimodal Composition Reasoning (MMStar [mmstar], BLINK [blink], ScienceQA [lu2022learn], M3CoT [chen-etal-2024-m3cot]). Additional descriptive details can be found in the Appendix.
Baselines.
We evaluate our method against a comprehensive set of state-of-the-art baselines, which can be categorized into four paradigms: (1) Zero-shot VLMs (GPT-4o [gpt4v], LLaVA-OneVision [li2024llava], InternVL3.5-8B [wang2025internvl35], Qwen2.5-VL-7B [bai2025qwen25vl]), (2) Explicit CoT-based methods (SCAFFOLD [lei2024scaffolding], ICoT [gao2024interleaved], Multimodal-CoT [zhang2023multimodal], CCoT [MitraCCoT], Chain-of-Focus [zhang2025chain]), (3) Visual Enhanced methods, including tool-augmented reasoning (DeepEyes [zheng2025deepeyes]) and RL-enhanced VLM reasoning (Vision-R1 [huang2025vision], PAPO [wang2025perception], VL-Rethinker [wang2025vl]), and (4) Multimodal Latent Reasoning approaches (Laser [li2025latent], LVR [li2025lvr], Monet [wang2025monet], DMRL [liu2025reasoningminddynamicmultimodal]).
Training Dataset.
To facilitate reproducibility, we first detail the training data configuration. During the supervised fine-tuning (SFT) stage, we curate a subset of approximately 30K samples from OneThinker [DBLP:journals/corr/abs-2512-03043], selected for its diverse distribution of reasoning chain lengths. The statistical distributions of CoT length, reasoning steps, and topic categories within this subset are illustrated in Fig. 3.
Backbone Models.
To comprehensively evaluate the effectiveness and scalability of our approach, we instantiate our method across a diverse set of backbone architectures, including Qwen2-VL-2B/7B [wang2024qwen2], Qwen2.5-VL-3B/7B [bai2025qwen25vl], and Chameleon-7B [team2024chameleon].
Implementation Details.
We employ DeepSpeed ZeRO-2 optimization without CPU offloading, with a per-GPU batch size of 8. Training utilizes the Adam optimizer with a learning rate of $4\times 10^{-5}$ and $\beta_{1}=0.9$. For the proposed self-distillation loss, we set the weighting coefficient to $\lambda=1.0$ for 2B/3B-scale models and $\lambda=0.2$ for 7B-scale models, respectively. Input images are resized to different resolutions according to training dataset, please refer to Appendix for detailed preprocessing protocols. All experiments are conducted on 16 NVIDIA H20 GPUs (96GB VRAM each).
4.2 Overall Quantitative Results
Reasoning Accuracy.
As summarized in Table 1, we train and evaluate our method on three datasets that provide fine-grained CoT annotations. With the same training protocols, our approach achieves the best performance across both model backbones and all three datasets. Compared to traditional explicit CoT methods, our method yields clear-cut improvements, with an average accuracy gain of nearly 30% built upon the Qwen-2-VL architecture. Our approach further surpasses the second-best latent-based reasoning baseline (IVT-LR) by an average of +2.63% across all benchmarks.
As shown in Table 2, we further evaluate our method against widely-used mainstream benchmarks that diagnose multi-faceted reasoning capabilities. On most benchmarks, our approach achieves consistent improvements. Particularly on vision-centric benchmarks, our method attains remarkable gains in overall scores compared to previous state-of-the-art methods. Notably, we observe the most substantial improvement on MMVP (+4.67%), which employs CLIP-blind patterns and emphasizes the requirement for fine-grained visual discrimination. We attribute these gains to our designed RCF-VR. By maintaining a set of visual latents with fine-grained spatially-coherent constraints, our method effectively mitigates hallucination while capturing fine-grained visual details. Conversely, our approach exhibits modest performance on text-intensive benchmarks, notably SeedBench-2-Plus. This gap likely arises from the scarcity of domain-specific training data, underscoring the potential benefits of targeted fine-tuning for such tasks. Remarkably, our method achieves competitive performance against computationally intensive alternatives, including explicit Chain-of-Thought (CoT) and tool-augmented frameworks. Despite operating purely within the latent space—without relying on external knowledge retrieval or reinforcement learning-based optimization—our framework effectively captures rich semantic representations through token-wise depth scaling.
(a) CoT Length Distribution
(b) Thinking Step Distribution
(c) Topic Distribution
[…]
Reasoning Efficiency.
[…]
4.3 Ablation Analysis
Effect of Designed SCF-VR and RDS.
Table 4: Performance comparison with different mapping networks of $\mathcal{F}_{\text{SR}}$ on four kinds of backbones.
M3CoT (Acc %)
Method Linear Proj. Vanilla Conv. Point-wise Conv. [DBLP:journals/corr/abs-1712-05245]
Qwen2-VL-2B 51.2 43.2 43.7
Qwen2-VL-7B 66.4 61.4 62.3
Qwen2.5-VL-3B 65.3 61.1 62.6
Qwen2.5-VL-7B 73.9 69.2 69.8

### 11 May 2026 v4

@@ 3.2 Spatially-Coherent Finer Visual Replay
As mentioned earlier, we empirically reveal the gradient disparities between visual and textual tokens throughout the learning dynamics. Specifically, visual tokens consistently exhibit substantially smaller gradient Frobenius norms compared to textual tokens and latent token, indicating that visual representations remain under-optimized despite their critical role in multimodal reasoning, which aligns with the conclusions drawn in prior studies [42, 64]. Motivated by these insights, we introduce the visual replay module to reinforce the engagement of visual cues via salient region detection, while enhancing fine-grained spatially-coherent perception capabilities via self-distillation supervision at each reasoning step.
Attention-Guided Region Focus. Several works [32, 78] have demonstrated that LLMs exhibit fundamental visual grounding capabilities. To identify visually salient regions, we aggregate attention weights across all transformer layers and attention heads to obtain a consolidated spatial focus map. Specifically, given the $l$-th layer and the $h$-th attention head, we compute the mean attention map $\bar{\mathbf{A}}^{(t)}$ at reasoning step $t$:
@@
Spatially-Coherent Regularization. Although leveraging learned attention within Transformers provides explainable visual locations, it often suffers from limited spatial continuity due to the scattered nature of selected tokens and introduces noise associated with the attention sink phenomenon [66]. To mitigate these issues and enhance fine-grained perception without external annotations, we introduce self-distillation supervision. This mechanism involves cropping the visual regions exhibiting spatial coherence, re-encoding them, and supervising the visual latents with these high-fidelity features.
@@
where the density function $\mathcal{N}(r,c)$ counts the visited tokens falling within the window $\mathcal{N}(r,c)=\sum_{i\in\mathcal{I}^{(t)}}\mathbb{I}\big[r\leq r_{i}<r+W,c\leq c_{i}<c+W\big].$ Here, $(r_{i},c_{i})$ denotes the row-column position of token $i$ on the $G\times G$ grid, and $\mathbb{I}[\cdot]$ is the indicator function. This greedy selection ensures the cropped region captures a coherent visual context rather than scattered details. Then, the selected window is projected to pixel coordinates in the original image via mathematical transformation, cropped, and resized to the standard encoder input resolution using bilinear interpolation. We then re-encode this refined patch through the same visual encoder to obtain fine-grained representations $\{\mathbf{f}_{i}\}_{i=1}^{N_{v}^{\prime}}$. A global pooling operation $\mathrm{Pool}(\cdot)$ is applied to yield a robust reference token $\mathbf{u}^{\text{ref}}$. Finally, we align the coarse-grained global token $\mathbf{b}^{(t)}=\mathrm{Pool}(\mathbf{B}^{(t)})$ with this high-fidelity reference using a lightweight token super-resolution module $\mathcal{F}_{\text{SR}}:\mathbb{R}^{D}\to\mathbb{R}^{D}$. We minimize the reconstruction error as follows:
@@ 3.3 Routing Depth Scaling
While visual replay mechanisms significantly enhance fine-grained grounding by incorporating spatially-coherent visual latents, existing methods typically allocate uniform computational budgets across prefilling tokens and newly generated latents during contextual refinement. Intuitively, however, tokens and latents contribute heterogeneously to the final reasoning objective, rendering static budget allocation suboptimal. We hypothesize that a token’s processing requirement is intrinsically linked to its relative difficulty, which can be quantified via optimization dynamics.
A Theoretical Perspective: Rank Dynamics as a Proxy for Difficulty. To formalize difficulty, we analyze the optimization dynamics of QKVO weight matrices, positing that their evolution reflects sample complexity. For simple samples, consistent gradient updates drive weights toward a compact, low-rank structure where principal singular values dominate and noise is suppressed. Conversely, difficult samples involve conflicting contexts that force representations to constantly adjust between competing patterns. This instability manifests as resurrected suppressed singular values or reshaped principal distributions, causing persistent fluctuation in the nuclear norm of matrix. We quantify these dynamics via the nuclear norm, defined as follows:
Definition 3.1 (Nuclear Norm of a Matrix).
Let ${\bm{A}}\in\mathbb{R}^{m\times n}$ be a weight matrix in transformer, the Singular Value Decomposition (SVD) of $\bm{A}$} is given by ${\bm{A}}={\bm{U}}{\bm{S}}{\bm{V}}^{\top}$, where ${\bm{U}}\in\mathbb{R}^{m\times m}$ and ${\bm{V}}\in\mathbb{R}^{n\times n}$ are orthogonal matrices, and ${\bm{S}}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the non-negative singular values $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{r}>0$, with $r=\mathrm{rank}({\bm{A}})$. The nuclear norm of $\bm{A}$}, denoted as $\|{\bm{A}}\|_{*}$, is defined as the sum of its singular values:
$\|{\bm{A}}\|_{*}=\sum_{i=1}^{\min(m,n)}\sigma_{i}({\bm{A}})=\mathrm{Tr}\left(\sqrt{{\bm{A}}^{\top}{\bm{A}}}\right).$
Based on above defined difficulty indicator, we reveal heterogeneous optimization complexities across tokens during latent training, termed the fixed-depth optimization dilemma. To address this without modifying the pretrained VLM architecture while effectively leveraging its inherent knowledge, we introduce a lightweight router that dynamically allocates additional reasoning steps exclusively to critical tokens, enabling adaptive reasoning depths.
Router Network. Let the input token sequence at the $t$-th reasoning step be denoted as $\mathbf{U}^{(t)}=[\mathbf{X}\|\mathbf{B}^{(1)}\|\mathbf{z}^{(1)}\|\cdots\|\mathbf{B}^{(t)}\|\mathbf{z}^{(t)}]\in\mathbb{R}^{P^{(t)}\times d}$, where $P^{(t)}$ represents the total sequence length and $d$ is the hidden dimension. To facilitate the illustration of our depth scaling mechanism, we decompose the forward pass into layer-wise computations. Specifically, within the $l$-th transformer layer, we first compute the intermediate feature $\mathbf{H}^{(l,t)}=f(\mathbf{U}^{(t)})\in\mathbb{R}^{P^{(t)}\times d}$, where $f(\cdot)$ denotes the transformation of the $l$-th transformer layer in LLM. Subsequently, a lightweight router network computes a scalar importance score for each token based on its corresponding hidden representation:
@@
For notational brevity, we omit the layer index $l$ and reasoning step $t$, denoting refinement depth by $d$ (e.g., $\mathbf{s}^{(l,t)}\to\mathbf{s}^{(d)}$). Let $\mathbf{h}^{(0)}_{i}$ denote the initial hidden state of token $i$, and $\mathbf{m}^{(d)}$ the attention mask at step $d$. A router network dynamically selects a critical token subset $T_{\alpha}(\mathbf{s}^{(d)})$ based on importance scores, acting as a binary gate. Consequently, the transformation $f(\cdot)$ is applied exclusively to tokens $i\in T_{\alpha}(\mathbf{s}^{(d)})$ for $d$ refinement steps, while non-critical tokens retain their prior states to ensure efficiency. Finally, we aggregate the refined representations with step-aware positional encodings to incrementally form the final hidden state:
@@
This adaptive strategy allocates deeper reasoning pathways to critical tokens, enabling the model to capture complex visual contexts and facilitate more profound reasoning capabilities. We further provide a theoretical analysis demonstrating that, under mild assumptions, our design achieves a tighter generalization bound (see Appendix).
Theorem 3.1 (Generalization Analysis of Token-wise Depth Scaling).
Let $\mathcal{T}_{\text{std}}$ denote the standard $L$-layer Transformer with total parameter count $W_{\text{std}}=\Theta(Ld^{2})$. Let $\mathcal{T}_{\text{topk}}$ denote the token-wise depth scaling variant with Top-K selection, having total parameter count $W_{\text{topk}}=W_{\text{std}}+Md=W_{\text{std}}(1+O(1/d))$, where $M=\lceil L/2\rceil$ is the number of modified layers. Under Assumptions A.1A.6, with probability at least $1-\delta$, the population risk satisfies:
$\mathcal{L}_{\mathcal{D}}(f_{\text{topk}})\leq\mathcal{L}_{\mathcal{D}}(f_{\text{std}})-\frac{\delta_{0}KL}{4n}\cdot\eta_{\text{eff}}+\tilde{\mathcal{O}}\left(\sqrt{\frac{dL\log(nL/\delta)}{N}}\right)-\Omega\left(\frac{K^{2}L}{n^{2}}\sqrt{\frac{W_{\text{std}}}{N}}\right),$ (12)
where $\delta_{0}>0$ is the minimal improvement from single-layer iteration, $K=\rho n$ is the selection number, $\eta_{\text{eff}}$ is the effective selection rate, $d$ is the hidden dimension, and $N$ is the number of training samples.
3.4 Training Procedure
Curriculum Latent Training. To reduce annotation overhead and avoid human priors limiting learning, we do not directly supervise intermediate latent representations. Instead, we introduce a curriculum that links implicit latents with explicit reasoning chains. We first train the model with standard Chain-of-Thought (CoT) supervision, generating all reasoning steps explicitly to build baseline reasoning ability. Then, over training, we gradually replace one explicit step at a time with an informative $\langle\text{latent}\rangle$ token. This curriculum enables each latent to progressively ground relevant contextual cues, internalizing explicit reasoning into compact latent representations.
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda\frac{1}{T_{r}}\sum_{t=1}^{T_{r}}\mathcal{L}_{\text{recon}}^{(t)},~~~\mathcal{L}_{\text{CE}}=-\sum_{t=1}^{T_{r}+T_{a}}\log p_{\theta}\left(a_{t}|\mathbf{U}^{(t)},a_{<t}\right),$ (13)
where $\mathcal{L}_{\text{CE}}$ is the standard cross-entropy loss over the language modeling task, where $\lambda$ is a hyperparameter balancing the primary task and the auxiliary reconstruction objective. During inference, only the LLM component is active, and the visual play module is disabled, ensuring no additional computational overhead at test time.
Table 1: Comparison of various multimodal reasoning baselines across three benchmarks. We selected three datasets featuring detailed reasoning chains for training and evaluate on their test split, respectively. Three metrics are reported: Accuracy (%), Average number of Autoregressive Steps (# AR steps), and Average Generation Time (AVG. Time). denotes the reimplementation for the methods with the same configuration with ours. No-CoT notes that directly predicts answers without generating intermediate steps.
@@ 3.4 Training Procedure
Multimodal CoT [80] 42.5 106.3 3.10 58.3 83.9 2.44
CCoT [39] 44.1 177.2 5.31 63.8 164.0 5.23 51.2 76.4 7.21
ICoT [13] 46.0 96.5 2.86 65.4 77.4 2.28
SCAFFOLD [26] 44.9 170.8 5.14 62.5 162.3 4.91 48.7 72.8 6.72
Chain-of-Focus [79] 64.3 185.7 2.63 91.2 162.3 2.09 61.8 128.6 3.01
@@
Multimodal CoT [80] 30.6 110.5 3.62 50.7 98.7 3.33
CCoT [39] 31.4 168.4 5.35 51.3 174.2 5.39 33.1 150.6 5.31
ICoT [13] 32.3 110.9 5.43 53.4 92.4 4.62
SCAFFOLD [26] 31.1 194.3 6.12 47.5 160.6 6.03 32.8 156.0 4.17
Chain-of-Focus [79] 36.5 739.4 3.09 61.2 717.1 2.56 34.6 360.4 2.98
@@
Table 2: Performance comparison of our method against baselines across four paradigms. Benchmarks are categorized into three domains: Visual Perception, Compositional Reasoning, and Mathematical Reasoning. Among latent reasoning approaches, the best results are highlighted in bold, and the second-best are underlined. Our method is built upon the Qwen2.5-VL-7B backbone.
@@
GPT-4o [41] 68.70 72.00 68.00 64.70 63.80 30.39 31.80
Qwen2.5-VL-7B [2] 65.67 65.31 56.57 68.25 53.60 59.70 68.20 25.60 37.50
LLaVA-OneVision [29] 9M 74.00 61.22 51.10 63.00 49.34 59.13 58.60 - -
InternVL3.5-8B [60] 70K 57.67 69.78 56.15 59.38 54.81 53.33 71.60 28.30 -
@@
Multimodal CoT [80] 68.10 54.11 63.60 - - 57.90 56.40 21.80 35.60
CCoT [39] 69.00 68.95 64.90 - - 58.70 57.80 22.50 36.30
ICoT [13] 69.30 70.27 65.50 - - 60.40 58.90 23.30 37.00
@@
PAPO [63] 39K 68.67 54.11 57.52 68.12 52.66 45.80 67.53 - -
Vision-R1 [24] 200K 72.67 68.95 63.83 75.12 52.71 62.67 52.40 - 40.20
VL-Rethinker [55] 39K 72.67 70.27 71.08 63.50 55.55 63.20 72.80 29.30 -
DeepEyes [83] 47K 70.00 69.08 62.57 69.12 51.08 58.73 70.10 26.60 -
@@
LVR [27] 470K 64.00 47.39 65.19 53.62 53.60 57.93 - - -
Monet [58] 125K 68.00 65.88 56.36 68.00 50.71 60.33 - - -
DMLR [33] 70.10 - 65.80 - 56.92 60.27 59.10 24.40 38.80
Laser [62] 267K 72.00 70.05 67.72 72.50 - 60.10 - - -
@@
Table 3: We perform comprehensive ablation studies to analyze the contribution of each design choice. Experiments are conducted on three CoT benchmarks with multiple backbone models. To save the training overhead, all models are trained for 16 epochs with a balanced sampled subset.
@@
Table 4: Ablation Experiments. We provide ablation analysis of key parameters and experimental settings on three benchmarks. All the variants adopt Qwen2.5-VL-3B as the base model. Base refers the model that adopts RDS and SCF-VR.